please help.. they are indirect and direct variations, please show work.. asap
3.If y varies directly with x, and y = 8 when x = 16, find y when x = 46.
5. If y varies directly with x, and y = 2 when x = 4, find y when x = 45.
i got this k=2/4=1/2(45)y=90
7.If y varies indirectly with x, and y = 10 when x = 25, find y when x = 35.
i got k=10/25=2/5(35)y=14
some one please help me i really need to understand these problems they are due tonight. thanks a bunch if some one will help me//

Answer by RAY100(1637)

(3)y= kx,,,,,,,that is y varies directly with x
.
8 = k * (16),,,,,,,when y=8, x=16,,,,,,find k
.
k=8/16 =1/2
.
Variation is,,,,y = 1/2 * x
.
when x= 46,,,,,,,,y = (1/2) * 46 = 23
.
check,,,8/16 = 23/46,,,,,ok
.
(5) y=kx,,y varies directly with x
.
2=k*4,,,,when y=2,,,x=4,,,,find k
.
k=1/2
.
Variation is ,,,,y= 1/2 *x
.
y= (1/2)*45 =22.5
.
check 22.5 /45 = 2/4,,,,ok
.
(7) y= k/x,,,,,,,y varies in directly with x
.
10=k/25, y=10 when x=25,,,find k
.
250 = k
.
Variation is,,,,y = 250/x
.
y= 250/35 = 7.14,,,,,find y if x=35
.
check,,,y=k/x,,,xy =k = 10 * 25 = 7.14 * 35 = 250,,,ok
.